a) Give parametric equation (point-direction form) of the line which lies on both of the planes:

x + y + z = 1 and d x + 2y + 10z = 2. What is the direction d of this line?

b) Let n1 and n2 be the normal vectors to the two given planes. Without actual computation,

describe the relationship between d and n1 × n2.


1
Expert's answer
2021-10-05T13:42:59-0400

a) if Ax + Bx +Cz =D is the equation of a plane , then the vector normal to the given plane is "n_1 =i + 2 j -k"

Equation of f2 is "3x-y+4z=2"

Therefore the normal vector "n_2= 3i - j +4k"


b) We have "n_1 =i + 2 j -k" and "n_2= 3i - j +4k."

Then "n_1" and "n_2" are perpendicular to the line of intersection. Therefore "v = n_1 * n_2" is the perpendicular to both "n_1" and "n_2" and hence parallel to the plane line of intersection

"\begin{bmatrix}
 i & j & k \\
 1 & 2 & -1\\
 3 & -1 & 4\\
\end{bmatrix}= i(8-1)-j(4+3)+k(-1-6)=7i-7j-7k"